This question has two parts (10 pts): (a) Consider the freezing point of 0.01M KCl(aq), tf(KCl), compared with that of 0.01 M Ca(NO3)2(aq), tf(Ca(NO3)2. Which of the following is true (circle one only)? i. tf(KCl) > tf(Ca(NO3)2) ii. tf(KCl) < tf(Ca(NO3)2) iii. tf(KCl) = tf(Ca(NO3)2) (b) What volume (in mL) of phenol (Mm = 94.1 g/mol; nonelectrolyte; density = 1.07 g/mL) must be added to 25.0 g of benzene(l) to have the freezing point of the resulting solution be ?0.80ºC? The kf (benzene) = 5.1 ºC/m. The freezing point of pure benzene is 5.5ºC.
Answer #1

a)

KCl dissociates as

KCl \rightarrow K+ + Cl- .

Hence, number of ions = 2.

As it is completely dissociates, then vant hoff factor (i) = 2.

Again Ca(NO3)2 being strong electrolyte dissociates as

Ca(NO3)2\rightarrow Ca2+ + 2NO3-

So, number of ions = 3, then i = 3.

Now, freezing point depression;  \DeltaTf = i* Kf m.

So, 0.01 M Ca(NO3)​​​​​​2 has higher value in depression of freezing point than 0.01 M KCl.

Therefore, 0.01 M Ca(NO3)2 has lower freezing point than 0.01 M KCl.

So, tf(KCl)> tf(CaNO3)2 .

Option i is correct.

B)

By Roult's law for nonelectrolyte

\DeltaTf = Kf*m

Or, \DeltaTf = Kf * W2* 1000 MW LI

W2 is mass of solute (phenol)

M is molar mass of solute = 94.1 g/mol.

W1 is mass of solvent ( benzene) = 25.0 g.

Now, given freezing point depression = \Delta Tf  = (Tf - Tf' )

= (5.5 - 0.80) = 4.70c .

4.7 = 5.1 W2 1000 94.1 25.0

Or, W2 = (4.7*94.1*25.0)/(5.1*1000) =  2.168 g.

Given, Density of phenol = 1.07 g/mL.

Then , Volume of phenol = (2.168/1.07) = 2.026 mL.